OpenStudy (anonymous):

The Nutty Shoppe has 10 kg of mixed pecans and cashews which sell for $8.40 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of cashews are there in the mix?

OpenStudy (anonymous):



OpenStudy (anonymous):

work out the left side first

OpenStudy (anonymous):

that is pretty confuseing to me

OpenStudy (anonymous):

ok then 10 x 8.40

OpenStudy (anonymous):



OpenStudy (anonymous):

my computer wouldnt let me on the site

OpenStudy (anonymous):

The bag is 10kg, so c + p = 10. The total cost of the bag is $84.00, so 8c + 9p = 84. But you know c = 10-p from equation 1, above. Make that substitution into equation 2. Now: 8(10-p) + 9p = 84. 80 - 8p + 9p = 84 80 + p = 84 p = 4 Since c + p = 10, c must be 6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i get it now thanks bro

OpenStudy (anonymous):

don't forget to close the question